2.2. Solve the ivp sin(x) dx + y dy = 0, where y(0)


1
Expert's answer
2022-06-27T14:33:47-0400
"\sin(x) dx + y dy = 0, \quad y(0)=-2."

"y dy = -\sin(x) dx"

"y^2/2=\cos(x)+C"

"2=\cos(0)+C=1+C"

"C=1"

Finally

"y^2=2\cos(x)+2"

"y=\pm\sqrt{2\cos(x)+2}"

The initial condition requres negative value of "y" for specified "x", thus

"y=-\sqrt{2\cos(x)+2}"